OpenStudy (anonymous):

Got a test tomorrow help me out :)? How would I graph y=1/7|x+5|+2

OpenStudy (anonymous):

then compare it to tgraph of y=|x|

OpenStudy (anonymous):

please help!

OpenStudy (anonymous):

the 5 in the absolute value shifts the graph to the left by 5, the +2 shifts the equation up by 2 the 1/7 "flattens" the equation by a factor of 1/7 (for example, if x=1, abs(x) = 1, but (1/7)(abs(x))=1/7

OpenStudy (anonymous):

could oyu do a visual? still abit lost

OpenStudy (anonymous):

please.

OpenStudy (anonymous):

 just remember to plug in x values and solve for y. Since you are dividing by 7, finding x values that would make abs(x+5) = a value that is divisible by 7 (like 7 or 14) will make it easier to solve. (sorry for the poor graph...it's late here...) if it's still unclear, let me know

OpenStudy (anonymous):

thanks you first of all.